### v2

@@ 3.2 Service Contacted by Devices
First, regardless of the type of device, first party contacts constituted a significant proportion of all device communications. This indicates that the privacy policies of Chinese devices are relatively well-established. Second, the interaction with the support party is closely related to the device’s computational resource needs. Among the 12 device types, the camera, doorbell, speaker, and mirror interact more frequently with the support party. As a result, many manufacturers choose to collaborate with cloud service providers to ensure the devices can leverage cloud computing and storage capabilities, thereby enhancing their stability. The support parties contacted by these devices typically provide cloud computing and storage services (e.g., oss-cn-beijing.aliyuncs.com, ap-beijing.myqcloud.com), video streaming services (e.g., video.qq.com, stream.qqmusic.qq.com), and network communication services (e.g., dns.alidns.com).
@@
Compared with first and support parties, third party domains primarily involve data sharing, advertising services, geolocation, and personalized recommendation services. For example, the Huawei speaker contacts statist.tingmall.com for music streaming services, which may be related to third party audio components used by the device; the IMOU camera contacts api.weathercn.com for weather query services; the Huawei TV contacts several audio and video service providers, however, we did not find these service providers in their privacy policies. Additionally, we also observe that many devices access seemingly unrelated third party domains. For instance, Xiaomi speaker, Konka camera, Lenovo camera, Pisen camera, and Jooao camera all access the domain www.baidu.com. These observations suggest that while third party services offer various personalized features, they are not explicitly mentioned in privacy policies. As a result, unexpected destination contacts may raise user privacy concerns. Some third party service providers possess strong cross-platform data aggregation capabilities. They can integrate the user behaviors on different devices to build detailed user behavior profiles, which brings serious privacy risks [35, 37]. Therefore, device manufacturers should clearly define data usage boundaries when integrating third party services, transparently disclose relevant third party domains and their purposes in privacy policies, and allow users to make choices.
@@
Additionally, for 10 devices support firmware updates, there was no significant change in the number of third party service providers contacted. Overall, traffic destinations showed little change after the firmware updates, indicating that device firmware updates have little impact on traffic destinations in the short term.
@@
Commonly Contacted Organizations. To further assess whether the devices’ communications exhibit excessive reliance on specific organizations, we analyzed the organization of traffic destinations and ranked them based on interaction frequency. We focused on the companies most frequently contacted by consumer IoT devices and listed the top ten companies in Table 2. According to the table, 20 devices in the dataset contacted Google to use its DNS service. In contrast, the local DNS provider Greatbit, is used by only 17 devices, which is fewer than Google.
@@
In real-time communication for consumer IoT devices, stable and efficient DNS services are essential. While Google Public DNS is commonly used due to its global network and high availability, GreatBit, as a domestic provider, offers lower latency and better stability for local devices. However, the analysis shows that Google DNS is used more frequently than expected. In consumer IoT device real-time communication scenarios, stable and efficient DNS services are crucial. As a domestic service provider, GreatBit offers lower latency and higher stability for local devices, making it more suitable to meet the demands of domestic devices. Nevertheless, the analysis shows that compared with the local DNS service, the frequency of the device using Google DNS is higher. In addition to Google and Greatbit, the Organizations most frequently contacted by devices include Alibaba Cloud, Xiaomi, and Tencent. The services provided by these companies through their domains include cloud storage services (e.g., oss-cn-shenzhen.aliyuncs.com), NTP services (e.g., time1.cloud.tencent.com, ntp.aliyun.com), and video streaming and media content (e.g., findermp.video.qq.com, vv.video.qq.com), indicating that devices rely on these companies for the computational resources they provide.
@@ 3.3 Summary
By analyzing traffic destinations, we find that most Chinese devices rely on domestic services. Specifically, 99.75% of the traffic is directed to domestic servers that depend on domestic organizations. This means that Chinese IoT devices feature a high degree of localization and greater autonomy in control. From the perspective of lifecycle phases, we can infer that the main functionalities of IoT devices, e.g., controlling and monitoring, are built on domestic services. Most foreign services used by Chinese consumer IoT devices are basic Internet services like NTP and DNS, which can also be replaced by Chinese service providers. In addition to contacting first parties, many devices from different manufacturers commonly communicate with major domestic organizations such as Xiaomi, Alibaba, and Tencent. This allows them to receive information from many devices and potentially obtain the capability to construct user profiles, e.g., inferring types of devices in a household.
@@ 4.1 Encrypted Traffic Proportion Analysis
Consumer IoT device communications may involve sensitive information such as user personal data and surveillance audio or video. Transmitting such data in plaintext raises interception, tampering, and unauthorized risks. Thus, we leverage the entropy of the traffic to identify and analyze the proportion of unencrypted communications.
Common traffic analysis tools, such as Wireshark, can easily distinguish between encrypted protocols like TLS and QUIC, but cannot clearly classify customized encryption protocols. Therefore, we adapted the method proposed by Ren et al. [5] to determine whether traffic is encrypted by calculating the payload entropy of the packets. First, for HTTP protocol packets, we classify them into text, compressed, or media based on the context_type field. For the SSL protocol, if the payload entropy is greater than 0.8, it is classified as encrypted. For the DNS protocol, packets containing the dns_key field are treated as encrypted. For packets that cannot be classified in the first step, we further classify them based on the magic_number of the payload. The magic_number refers to a unique identifier or pattern of bytes in a packet that helps to determine its type or format, often used to identify specific file types or data structures. If the packet is identified as compressed or media data, it is categorized as compressed or media. Finally, for the remaining packets, we perform a final classification based on entropy values: packets with entropy greater than 0.9 are classified as encrypted, those with entropy less than 0.4 are classified as text, and those with entropy between 0.4 and 0.9 are classified as unknown. Ultimately, we classify the traffic into encrypted traffic, unencrypted traffic, and unknown traffic.
Figure 3 presents a heat map visualization illustrating the proportional distribution of encrypted, unknown, and unencrypted traffic across various lifecycle phases of consumer IoT devices.
@@
The heatmaps show that the proportion of unencrypted traffic is generally low, with no device type exceeding 10% of unencrypted traffic during the full device lifecycle. This indicates that encrypted communication dominates the data transmission of most devices. However, significant variation exists across device types. For instance, camera, doorbell, speaker, and TV devices exhibit an average encrypted traffic share of over 60%, largely driven by the security and privacy requirements in audio and video data transmission. In contrast, the humidifier, pet feeder, and plug devices have lower average encrypted traffic proportions of 36%, 34%, and 38%, while their unknown traffic proportions are the highest among all devices, at 50%, 54%, and 49%, respectively. Analysis of traffic entropy reveals that most of these devices’ traffic falls within the entropy range of 0.4-0.7. Upon further inspection, it appears that these devices do not use standard encryption protocols but instead rely on proprietary protocols developed by the manufacturers. These proprietary protocols have unstable entropy. By reverse engineering, we discovered that some devices can deduce the syntax and semantics of the manufacturer’s protocols. Therefore, we suggest that manufacturers can adopt standardized encryption protocols, such as TLS 1.3. In cases where proprietary protocols must be used, they can be combined with packet padding and traffic obfuscation techniques [36] to improve resistance against protocol reverse engineering and effectively reduce the risks of data leakage.
@@ 4.2 Encryption Protocol Analysis
Through analysis, we found that TLS 1.2 is widely used for encrypted communication across devices, with 62 devices using TLS 1.2 at least once. However, some devices (e.g., IMOU doorbells, Pisen cameras, and Tenda cameras) continue to rely on insecure SSL or TLS 1.1 protocols, which accounts for 7.6% of the total communication traffic on average. The devices generally use these protocols during the setup phase and the interaction phase. The setup and interaction phases often involve the transmission of sensitive data such as user identity information, remote control commands, and status update information. If weak encryption protocols are used during these phases, attackers may exploit them through MITM attacks to intercept and reconstruct the transmitted data. Currently, protocols below TLS 1.2 have been deemed insecure by multiple international standards organizations. Devices that continue to use such protocols may fail to comply with data protection requirements such as the General Data Protection Regulation (GDPR) and China’s Cybersecurity Law. Therefore, device manufacturers should review and phase out insecure protocols such as SSL, TLS 1.0, and TLS 1.1, and adopt TLS 1.2 or higher as the standard for secure communication.
Analyzing the destination IP addresses of this traffic showed that most targets are telecommunications service providers (e.g., China Mobile) or cloud service providers (e.g., Alibaba Cloud). Moreover, not all encrypted communications from these devices use SSL or TLS 1.1. There is a mixed use with other protocols (e.g., TLS 1.2). This may result from manufacturers neglecting to fully deprecate outdated protocols during certain functionality upgrades, leaving some connections still using SSL or TLS 1.1. On the other hand, while TLS 1.3 offers significant improvements in performance and security, only three devices in our dataset use it. This indicates that despite its clear advantages, its adoption remains limited, and manufacturers must make greater efforts to implement the latest security protocols. Additionally, we identified 10 devices (e.g., Aqara Camera, Aqara Plug, Xiaovv Camera) using proprietary protocols developed by the manufacturers. The security of these protocols remains to be further analyzed.
@@
Additionally, for the 10 devices that support firmware updates, most used TLS 1.2 for encryption before the update, while three devices relied on proprietary protocols developed by the manufacturers. The XiaobaiY2 camera also used SSL in addition to TLS 1.2. After the update, all devices upgraded their encryption protocols. In addition to continuing to use TLS 1.2, they also adopted TLS 1.3. The three devices that previously used proprietary protocols switched to TLS 1.2, and the XiaobaiY2 camera stopped using SSL, now relying only on TLS 1.2 and TLS 1.3 for encrypted communication.
@@ 4.3 TLS Certificate Security Analysis
The TLS 1.2 protocol is the most widely used encryption protocol in our dataset. To further understand its security, this section presents a statistical analysis of the certificate for the 63 devices in the dataset that use the TLS 1.2 protocol. The results show that 48 devices use a public key certificate model for communication. The remaining 15 devices used the Pre-Shared Key (PSK) model without relying on certificates. Given the crucial role digital certificates play in TLS encryption security, we further analyzed the self-signed certificate usage and the encryption algorithms’ security. Additionally, we conducted dynamic tests using the mitmproxy to assess the validity of certificates when devices powered on again.
@@ 4.3.1 Certificate Usage
Overall, the certificate security of domestic devices is generally good, but some insecure certificates were still identified. Table 4 statistics the number of TLS implementations that may be insecure, including insecure signature algorithms, insecure cryptographic algorithms, and self-signed certificates. Among the 48 devices, 2 use insecure algorithms, such as the SHA1 signature algorithm and the RSA-1024 cryptographic algorithm, primarily for communication with Aliyun and AWS servers. These algorithms are now considered insecure and vulnerable to attacks. Therefore, device manufacturers can regularly audit the cypher suites and certificate configurations used in their products, eliminate algorithms that are widely recognized as insecure, and ensure that their TLS implementations comply with current industry security standards.
@@ 4.3.2 Certificate Validation
To investigate this issue, we first successfully captured traffic during device reconnections and discovered that 40 devices would conduct TLS connections with certificate exchange with 106 servers during the reconnection period. Next, we conducted a mitmproxy certificate replacement attack on these TLS connections. The results are presented in Table 5.
@@
The results show that approximately 65% of the servers responded with errors, such as ”Unknown CA,” ”Decrypt error,” ”Close notify” alerts, or actively terminated the connection, and 35% of the servers associated with 23 devices did not verify the certificates. We decrypted the traffic between the devices and servers and extracted APIs used by the devices, revealing critical device information. For example, the API “/S_PROD1_MEIDIMINI/…/215413.log” exposes device log data; the API “/v3/am/query_device_bind_status” reveals device binding status, and in this API, the fields “device_name” and “device_sk” are exposed. We suspect that “device_sk” may be related to the device’s secret key, which is typically used to verify the identity and ensure the security of the device. These APIs expose sensitive information, if exploited by attackers, they could lead to serious consequences such as log data leakage, unauthorized remote control of the device, or tampering with its binding status. These findings suggest that many consumer IoT devices adopt insecure certificate validation mechanisms, making them vulnerable to certificate replacement attacks. We have reported this issue to the respective device manufacturers and submitted the vulnerability to the China National Vulnerability Database (CNVD).
@@ 4.4 PII Exposure in Unencrypted Traffic
Personally Identifiable Information (PII) refers to any data that can identify an individual [16]. Common PII includes names, ID numbers, phone numbers, email addresses, etc. Unencrypted traffic containing PII significantly elevates the risk of user data exposure. In this section, we identify PII in unencrypted traffic using text matching and regular expressions to measure the security practice. To comprehensively analyze PII exposure, we input as much personal information as possible into the companion app when we collect traffic.

### v3

@@ 4.2 Servers contacted by devices
First, regardless of the type of device, first party contacts constituted a significant proportion of all device communications. On average, each device contacts 7.0 first party servers, 2.6 support party servers, and 2.2 third party servers. This indicates that the privacy policies of Chinese devices are relatively well-established. Second, the interaction with the support party is closely related to the device’s computational resource needs. Among the 12 device types, the camera, doorbell, speaker, and mirror interact more frequently with the support party. These devices, on average, contact 4.3 support party servers, while the remaining devices contact an average of 0.9 support party servers. Many manufacturers choose to collaborate with cloud service providers to ensure the devices can leverage cloud computing and storage capabilities, thereby enhancing their stability. The support party servers contacted by these devices typically provide cloud computing and storage services (e.g., oss-cn-beijing.aliyuncs.com, ap-beijing.myqcloud.com), video streaming services (e.g., video.qq.com, stream.qqmusic.qq.com), and network communication services (e.g., dns.alidns.com).
@@
Compared with first and support party servers, third party servers primarily involve data sharing, advertising services, geolocation, and personalized recommendation services. For example, the Huawei speaker contacts statist.tingmall.com for music streaming services, which may be related to third party audio components used by the device; the IMOU camera contacts api.weathercn.com for weather query services; the Huawei TV contacts several audio and video service providers, however, we did not find these service providers in their privacy policies. Additionally, we also observe that many devices access seemingly unrelated third party servers. For instance, Xiaomi speaker, Konka camera, Lenovo camera, Pisen camera, and Jooao camera all access the domain www.baidu.com. These observations suggest that while third party servers offer various personalized features, they are not explicitly mentioned in privacy policies. As a result, unexpected destination contacts may raise user privacy concerns. Some third party servers possess strong cross-platform data aggregation capabilities. They can integrate the user behaviors on different devices to build detailed user behavior profiles, which brings serious privacy risks [39, 41]. Therefore, device manufacturers should clearly define data usage boundaries when integrating third party services, transparently disclose relevant third party domains and their purposes in privacy policies, and allow users to make choices.
@@
Additionally, for 10 devices that support firmware updates, there was no significant change in the number of third party service providers contacted. Overall, traffic destinations showed little change after the firmware updates, indicating that device firmware updates have little impact on traffic destinations in the short term.
@@
Commonly contacted organizations. To further assess whether the devices’ communications exhibit excessive reliance on specific organizations, we analyzed the organization of traffic destinations and ranked them based on interaction frequency. We focused on the companies most frequently contacted by consumer IoT devices and listed the top ten companies in Table 3. According to the table, 20 devices in the dataset contacted Google. By analysing the traffic, we find that devices contact Google to use its DNS service. In contrast, the Chinese DNS provider Greatbit, is used by only 17 devices, which is fewer than Google.
@@
In consumer IoT device real-time communication scenarios, stable and efficient DNS services are crucial. As a domestic service provider, GreatBit offers lower latency and higher stability for local devices, making it more suitable to meet the demands of domestic devices. Nevertheless, the analysis shows that compared with the local DNS service, the frequency of the device using Google DNS is higher. In addition to Google and Greatbit, the Organizations most frequently contacted by devices include Alibaba Cloud, Xiaomi, and Tencent. The services provided by these companies through their domains include cloud storage services (e.g., oss-cn-shenzhen.aliyuncs.com), NTP services (e.g., time1.cloud.tencent.com, ntp.aliyun.com), and video streaming and media content (e.g., findermp.video.qq.com, vv.video.qq.com), indicating that devices rely on these companies for the computational resources they provide.
@@ 4.3 Summary
By analyzing traffic destinations, we find that most Chinese devices rely on domestic services. Specifically, 99.75% of the traffic is directed to domestic servers that depend on domestic organizations. This means that Chinese IoT devices feature a high degree of localization and greater autonomy in control. From the perspective of lifecycle phases, we can infer that the main functionalities of IoT devices, e.g., controlling and monitoring, are built on domestic services. Most overseas services used by Chinese consumer IoT devices are basic Internet services like NTP and DNS, which can also be replaced by Chinese service providers. In addition to contacting first parties, many devices from different manufacturers commonly communicate with major domestic organizations such as Xiaomi, Alibaba, and Tencent. This allows them to receive information from many devices and potentially obtain the capability to construct user profiles, e.g., inferring types of devices in a household.
@@ 5.1 Encrypted traffic proportion analysis
Consumer IoT device communications may involve sensitive information such as user personal data and surveillance video. Transmitting such data in plaintext raises interception, tampering, and unauthorized risks. Thus, we leverage the entropy of the traffic to identify and analyze the proportion of unencrypted communications.
Common traffic analysis tools, such as Wireshark, can easily distinguish encrypted protocols like TLS and QUIC, but cannot clearly classify customized encryption protocols. Therefore, we adapted the method proposed by Ren et al. [9] to determine whether traffic is encrypted by calculating the payload entropy of the packets. First, for HTTP protocol packets, we classify them into text, compressed, or media based on the context_type field. For the SSL protocol, if the payload entropy is greater than 0.8, it is classified as encrypted. For the DNS protocol, packets containing the dns_key field are treated as encrypted. For packets that cannot be classified in the first step, we further classify them based on the magic_number of the payload. The magic_number refers to a unique identifier or pattern of bytes in a packet that helps to determine its type or format, often used to identify specific file types or data structures. If the packet is identified as compressed or media data, it is categorized as compressed or media. Finally, for the remaining packets, we perform a final classification based on entropy values: packets with entropy greater than 0.9 are classified as encrypted, those with entropy less than 0.4 are classified as text, and those with entropy between 0.4 and 0.9 are classified as unknown. Ultimately, we classify the traffic into encrypted traffic, unencrypted traffic, and unknown traffic.
Figure 4 presents a heat map illustrating the proportional distribution of encrypted, unknown, and unencrypted traffic in various lifecycle phases of consumer IoT devices.
@@
The heatmaps show that the proportion of unencrypted traffic is generally low, with no device type exceeding 11% of unencrypted traffic in the full lifecycle. This indicates that encrypted communication dominates the data transmission of most devices. However, significant variation exists between different device types. For instance, camera, doorbell, speaker, and TV devices exhibit an average encrypted traffic share of over 60%, largely driven by the security and privacy requirements in audio and video data transmission. In contrast, the humidifier, pet feeder, and plug devices have lower average encrypted traffic proportions of 36%, 34%, and 38%, while their unknown traffic proportions are the highest among all devices, at 50%, 54%, and 49%, respectively. Analysis of traffic entropy reveals that most of these devices’ traffic falls within the entropy range of 0.4-0.7. Upon further inspection, it appears that these devices do not use standard encryption protocols but instead rely on proprietary protocols developed by the manufacturers. These proprietary protocols have unstable entropy. By reverse engineering, we discovered that some devices can deduce the syntax and semantics of the manufacturer’s protocols. Therefore, we suggest that manufacturers adopt standardized encryption protocols, such as TLS 1.3. In cases where proprietary protocols must be used, they can be combined with packet padding and traffic obfuscation techniques [40] to improve resistance against protocol reverse engineering and effectively reduce the risks of data leakage.
@@ 5.2 Encryption protocol analysis
Through analysis, we found that TLS 1.2 is widely used for encrypted communication of devices, with 62 devices using TLS 1.2 at least once. However, some devices (e.g., IMOU doorbell, Pisen camera, and Tenda camera) continue to rely on insecure SSL or TLS 1.1 protocols, which account for 7.6% of the total communication traffic on average. The devices generally use these protocols during the setup phase and the interaction phase. The setup and interaction phases often involve the transmission of sensitive data such as user identity information, remote control commands, and status update information. If weak encryption protocols are used during these phases, attackers may exploit them through MITM attacks to intercept and reconstruct the transmitted data. Currently, protocols below TLS 1.2 have been deemed insecure by multiple international standards organizations. Devices that continue to use such protocols may fail to comply with data protection requirements such as the General Data Protection Regulation (GDPR) and China’s Cybersecurity Law. Therefore, device manufacturers should review and phase out insecure protocols such as SSL, TLS 1.0, and TLS 1.1, and adopt TLS 1.2 or higher for secure communication.
Analyzing the destination IP addresses of the traffic showed that most targets are telecommunications service providers (e.g., China Mobile) or cloud service providers (e.g., Alibaba Cloud). Moreover, not all encrypted communications from these devices only use SSL or TLS 1.1. There is a mixed use with other protocols (e.g., TLS 1.2). This may result from manufacturers neglecting to fully deprecate outdated protocols during certain functionality upgrades, leaving some connections still using SSL or TLS 1.1. On the other hand, while TLS 1.3 offers significant improvements in performance and security, only five devices in our dataset use it. This indicates that despite its clear advantages, its adoption remains limited, and manufacturers must make greater efforts to implement the latest security protocols. Additionally, we identified 11 devices (e.g., Aqara Camera, Aqara Plug, Dogness Feeder) using proprietary protocols developed by the manufacturers. The security of these protocols remains to be further analyzed.
@@
Additionally, for the 10 devices that support firmware updates, most used TLS 1.2 for encryption before the update, while three devices relied on proprietary protocols developed by the manufacturers. The XiaobaiY2 camera also used SSL in addition to TLS 1.2. After the update, all devices upgraded their encryption protocols. In addition to continuing to use TLS 1.2, they also adopted TLS 1.3. The three devices that previously used proprietary protocols switched to use TLS 1.2, and the XiaobaiY2 camera stopped using SSL, now relying on TLS 1.2 and TLS 1.3 for encrypted communication.
@@ 5.3 TLS certificate security analysis
The TLS 1.2 protocol is the most widely used encryption protocol in our dataset. To further understand its security, this subsection presents a statistical analysis of the certificate for the 62 devices in the dataset that use the TLS 1.2 protocol. The results show that 48 devices use a public key certificate model for communication. The remaining 14 devices use the Pre-Shared Key (PSK) model without relying on certificates. Given the crucial role digital certificates play in TLS encryption security, we further analyzed the self-signed certificate usage and the security of the encryption algorithms. Additionally, we conducted dynamic tests by using mitmproxy to assess the validity of certificates when devices are powered on again.
@@ 5.3.1 Certificate usage
Overall, the certificate security of devices is generally good, but some insecure certificates were still identified. Table 5 statistics the number of TLS implementations that may be insecure, including insecure signature algorithms, insecure cryptographic algorithms, and self-signed certificates. Among the 48 devices, 2 use insecure algorithms, such as the SHA1 signature algorithm and the RSA-1024 cryptographic algorithm, primarily for communication with Aliyun and AWS servers. These algorithms are now considered insecure and vulnerable to attacks. Therefore, device manufacturers can regularly audit the cypher suites and certificate configurations used in their products, eliminate algorithms that are widely recognized as insecure, and ensure that their TLS implementations comply with current industry security standards.
@@
  • 4
    *Device lifecycle phases: SU (setup), ID (idle), IR (interaction), DE (deletion).
@@ 5.3.2 Certificate validation
To investigate this issue, we first captured the traffic during device reconnections. We discovered that 40 devices conduct TLS connections with certificate exchange with 106 servers during the reconnection period. Next, we conducted a mitmproxy certificate replacement attack on these TLS connections. The results are presented in Table 6.
@@
The results show that approximately 65% of the servers responded with errors, such as ”Unknown CA,” ”Decrypt error,” ”Close notify” alerts, or actively terminated the connection, and 35% of the servers associated with 23 devices communicate normally. We decrypted the traffic between the devices and servers and extracted APIs used by the devices, revealing critical device information. For example, the API “/S_PROD1_MEIDIMINI/…/215413.log” exposes device log data, the API “/v3/am/query_device_bind_
status” reveals device binding status, and in this API, the fields “device_name” and “device_sk” are exposed. We suspect that “device_sk” may be related to the device’s secret key, which is typically used to verify the identity and ensure the security of the device. These APIs expose sensitive information, if exploited by attackers, they could lead to serious consequences such as log data leakage, unauthorized remote control of the device, or tampering with its binding status. These findings suggest that many consumer IoT devices adopt insecure certificate validation mechanisms, making them vulnerable to certificate replacement attacks. We have reported this issue to the respective device manufacturers and submitted the vulnerability to the China National Vulnerability Database (CNVD).
@@ 5.4 PII exposure in unencrypted traffic
Personally Identifiable Information (PII) refers to any data that can identify an individual [20]. Common PII includes names, ID numbers, phone numbers, email addresses, etc. Unencrypted traffic containing PII significantly elevates the risk of user data exposure. In this subsection, we identify PII in unencrypted traffic using text matching and regular expressions to measure the security practice. To comprehensively analyze PII exposure, we input as much personal information as possible into the companion app when we collect traffic.